### v1

# Inflation with Gauss-Bonnet Correction and Higgs Potential

Abstract
We investigate the cosmological inflation for the Einstein-Hilbert action plus the Higgs potential function and the Gauss-Bonnet term coupled with the Higgs scalar field through a dilaton-like coupling. Then, using the Friedmann-Lemaître-Robertson-Walker metric and considering the appropriate slow-roll parameters, we derive the necessary equations of motion. In the proposed model, since the $e$-folding integral cannot be easily solved analytically, we first utilize a well-known Taylor expansion. Then, with a certain range of values derived for the model parameters, utilizing a mixture of several diagrams and numerical analysis methods, we obtain results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with observational data within the acceptable range of the $e$-folding values. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model.
PACS numbers: 98.80.Cq; 04.50.Kd; 98.80.Es; 98.80.-k
Keywords: Cosmological Inflation; Modified Gravity; Higgs Field; Gauss-Bonnet Term
1 Introduction
The cosmic inflation theory has been agreed upon by many cosmologists as a suitable way to explain fundamental problems of the standard model of cosmology, such as the flatness problem and the horizon problem in the early universe. For this purpose, the inflation generated based on a canonical scalar field with minimal coupling is one of the simplest methods available, wherein the scalar field slowly rolls during the inflation. In this type of model, the scalar field is called inflaton because it is the main cause of inflation [1, 2, 3, 4, 5, 6, 7]. According to this model, the quantum fluctuations of the inflaton scalar field affect the curvature perturbations, and with the growth of these perturbations, cosmic structures such as galaxies and stars are formed over time. To check the correctness of such models, the measurement of temperature anisotropy in the cosmic microwave background (CMB) can be used. The latest CMB data obtained from the Planck satellite show serious limitations on the values of the scalar spectral index of curvature perturbations ($n_{\rm S}$) and the tensor-to-scalar ratio ($r$). Such limitations indicate that there is not much freedom in choosing the parameters of the subject, including the type of inflaton and its potential [8, 9, 10].
In addition to such a simple model for inflation, other relatively complex models have also been presented by researchers to answer the issues and problems of this field. Although the inflation theory provides a consistent answer to the old problems of cosmology, despite the efforts made, new questions (along with ambiguities) have been raised, including the nature of the inflaton scalar field and the nature of dark matter and dark energy [11].
At present, the only scalar field available in the set of fundamental particles is the field corresponding to the Higgs boson, which appears to be a suitable candidate for the inflaton, see, e.g., Refs. [7, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In particular, with the discovery of Higgs features in the large hadron collider, which shows that the Higgs boson model has significant uncertainties in its parameters compared to some of the models used to develop the standard model. However, if the Higgs scalar field enters the inflation directly, it will be incompatible with the standard model of particle physics. To solve this problem, it has been proposed to add a non-minimal coupling to the gravity term, which has been studied the most, see, e.g., Refs. [12, 21, 19, 22]. By considering the non-minimal coupling between gravity and the Higgs scalar field, it is possible to remedy the aforementioned inconsistency problem and stabilize inflation, but new problems are encountered at the same time. Indeed, the price given for the desirability of this model is the unusual magnitude of the non-minimal coupling, approximately $10^{4}$. Such a (meta-)stability of the vacuum in the Higgs inflation model is still under discussion [23, 24]. On the other hand, due to the presence of the coupling term with a large coupling constant, the unitarity bound of the theory is violated because the energy scale in this inflation model is larger than or close to the cutoff value [22].
On the other hand, in the discussion of Higgs inflation, there is some freedom in the coupling of the Higgs scalar field with gravity, since measurements of the large hadron collider determine the coupling constant of the Higgs scalar field, but not its coupling constant with gravity. Hence, any arbitrary coupling constant with gravity can be used for it, see, e.g., Ref. [25]. In this work, we utilize this freedom of the Higgs scalar field and attempt to couple it with the Gauss-Bonnet term instead of the Ricci scalar.
The Gauss-Bonnet term is the second term of the Lovelock Lagrangian. It is also induced from the superstring theory and a subclass of the Horndeski theory. However, including the Gauss-Bonnet term alone in a four-dimensional Lagrangian of gravitation does not affect the equations of motion, see, e.g., Refs. [26, 27, 28]. Actually, it is topologically invariant in $4$-dimensions, and hence non-dynamical in $4$-dimensions. Nonetheless, its effect can appear when coupled non-minimally with a scalar field function, which can then play as an effective potential in causing inflation. Of course, it creates violent negative instabilities for tensor perturbations around a de Sitter background on small scales [9, 10, 11, 29, 30, 31].
The inflation resulting from the coupling with the Gauss-Bonnet term has been studied in Refs. [8, 9, 10, 11, 24, 25, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. For simplicity in calculations and analysis, in a number of references such as [8, 9, 54], the potential function and the coupling function ($V(\phi)$ and $\xi(\phi)$ as functions of the inflaton scalar field) are assumed so that the relaxation condition $V(\phi)\xi(\phi)$ is a constant. In these references, the focus is on two specific models, namely ($V(\phi)=V_{0}e^{-\lambda\phi}$ and $\xi(\phi)=\xi_{0}e^{\lambda\phi}$ with $V_{0}$ and $\xi_{0}$ as constants) and ($V(\phi)=V_{0}\phi^{n}$ and $\xi(\phi)=\xi_{0}\phi^{-n}$), wherein the results of calculations lead to the reduction of the tensor-to-scalar ratio.
In this study, we focus on breaking such a relaxation condition to check whether the obtained results are consistent with the Planck observational data. Moreover, we consider the inflaton scalar field to be a Higgs scalar field with a Higgs potential. However within inflation, since the field is much larger than the expected vacuum value, the potential can be approximated to a quartic form, $V_{0}\phi^{4}$. On the other hand, in this work, we consider the coupling function as dilaton-like, i.e., $\xi(\phi)=\xi_{0}e^{-\lambda\phi}$, and investigate the scalar spectral index and the tensor-to-scalar ratio, and compare the results with the most recent constraints by the Planck collaboration [55], i.e.,
$n_{\rm S}=0.9649\pm 0.0042\quad$
${}\qquad\left({\rm at}\ 68\%\,{\rm CL,PlanckTT,TE,EE+lowE+lensing}\right),$
$r<0.10\quad\left({\rm at}\ 95\%\,{\rm CL,PlanckTT+lowE+lensing}\right)$ (1)
and the joint Planck, BK1515 and BAO data in further constraint tightened the upper limit on r as
$r<0.056\quad{\rm at}\ 95\%\,{\rm CL}.$ (2)
Also, as another phenomenological aspect of the Gauss-Bonnet correction, we investigate the issue of gravitational waves, see, e.g., Ref. [56]. In this regard, by directly measuring the gravitational waves resulting from the merger of neutron star GW170817170817 [57] and comparing them to their electromagnetic counterpart GRB 170817170817[58], the observable range of the speed of gravitational waves, $c_{{}_{\rm GW}}(=c_{\rm T})$, has been determined to be
$3\times 10^{-15}\leq c_{{}_{\rm GW}}/c-1\leq 7\times 10^{-16}$ (3)
and equivalently
$6\times 10^{-15}<c_{{}_{\rm GW}}^{2}/c^{2}-1\leq 1.4\times 10^{-15},$ (4)
where $c$ denotes the speed of light and $c_{\rm T}$ shows the speed of tensor perturbation modes. As is evident, in reality, this order of difference is negligible to a certain extent.
In this regard, the work is organized as follows. In the next section, we first introduce the proposed action, which includes the Gauss-Bonnet term with a coupled scalar field. Then, we derive the modified Friedmann equations in the framework of Friedmann-Lemaître-Robertson-Walker (FLRW) and review the corresponding slow-roll inflation with the necessary equations. In Sec. III, we introduce the proposed model and proceed the calculations while considering the Higgs potential and a dilaton-like coupling function. In this section, we use a mixture of several plots and numerical analysis methods to get the best model parameters to determine acceptable inflationary observables. We also obtain the inflationary observables in the absence of the Gauss-Bonnet term. In addition, we consider the speed of gravitational waves for the proposed model during inflation. Finally, in Sec. IV, we summarize the concluding results obtained.
2 Slow-Roll Inflation with Gauss-Bonnet Correction
We consider the following action in a four-dimensional spacetime as [^1] Note that, the Higgs inflation action is usually considered by the coupling function $F(\phi)=1+f\phi^{2}$ (with $f$ as a constant) non-minimally coupled to the Ricci scalar in the Jordan frame. In this work, we consider $f=0$ without mentioning/resorting the Jordan frame. However, for the issue of the Jordan frame and/or the Einstein frame in this matter, see, e.g., Refs. [24, 54].
$S\!=\!\int\!d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}R-\dfrac{\omega}{2}\left(\nabla\phi\right)^{2}-V(\phi)-\frac{1}{2}\xi(\phi)R^{2}_{\rm GB}\right]\!,$ (5)
where $R$ is the Ricci scalar, $R^{2}_{\rm GB}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet term, and $\phi$ is the inflaton scalar field with a potential $V(\phi)$ and is non-minimally coupled to the Gauss-Bonnet term with the coupling function $\xi(\phi)$. In this study, we consider the coupling constant, $\omega$, to be $\omega=1$, and $\kappa\equiv 8\pi G/c^{4}$ also to be $\kappa=1$ in the Planckian units including the natural units, $\hbar=1=c$.
The variation of action (5) with respect to the inflaton scalar field gives the modified Klein-Gordon equation as
$\square\phi-V^{{}^{\prime}}(\phi)-\frac{1}{2}\xi^{{}^{\prime}}(\phi)R^{2}_{\rm GB}=0,$ (6)
where $\square={}_{;\alpha}{}^{;\alpha}$, the prime is the derivative with respect to $\phi$, and the lowercase Greek indices run from zero to three. Also, the variation of action (5) with respect to the metric yields the modified Einstein equation as
$(R^{\mu\nu}\!-\frac{1}{2}g^{\mu\nu}R)-\nabla^{\mu}\phi\nabla^{\nu}\phi+g^{\mu\nu}V(\phi)+\frac{1}{2}g^{\mu\nu}\nabla^{\lambda}\phi\nabla_{\lambda}\phi$ (7)
$-\!2g^{\mu\nu}R\,\square\xi+\!2R\nabla^{\mu}\nabla^{\nu}\xi-\!4\nabla_{\lambda}\nabla^{\mu}\xi R^{\lambda\nu}-\!4\nabla_{\lambda}\nabla^{\nu}\xi R^{\lambda\mu}$ (8)
$+4\,\square\xi R^{\mu\nu}+4g^{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\xi R^{\alpha\beta}-4\nabla_{\alpha}\nabla_{\beta}\xi R^{\mu\alpha\nu\beta}=0.$ (9)
Note that, only the terms related to the derivatives of $\xi(\phi)$ are appeared, wherein, according to the topological theorem, the terms with the coefficient $\xi(\phi)$ are omitted. In the continuation, we consider the spatially flat FLRW metric
$ds^{2}=-dt^{2}+a^{2}(t)\left(dx^{2}+dy^{2}+dz^{2}\right),$ (10)
where $a(t)$ is the scale factor as a function of the cosmic time $t$. Also, by accepting the homogeneity and isotropy, we consider the inflaton scalar field to be only a function of the cosmic time. Then, we insert metric (10) into Eqs. (6) and (7), and respectively get the first and second modified Friedmann equations and the modified Klein-Gordon equation as
$3H^{2}+\frac{1}{2}\dot{\phi}^{2}+V(\phi)+12H^{3}\dot{\xi}=0,$ (11)
$H^{2}\left(1-4\ddot{\xi}\right)-2\dfrac{\ddot{a}}{a}\left(1-4H\dot{\xi}\right)+V(\phi)-\frac{1}{2}\dot{\phi}^{2}=0,$ (12)
$12H^{2}\dfrac{\ddot{a}}{a}\xi^{{}^{\prime}}(\phi)+\ddot{\phi}+3H\dot{\phi}+V^{{}^{\prime}}(\phi)=0,$ (13)
where $H=\dot{a}/{a}$ is the Hubble parameter and in turn obviously $\ddot{a}/a=\dot{H}+H^{2}$, and the dot represents the time derivative. It is well-known that, knowing the functionality of $V(\phi)$ and $\xi(\phi)$ in term of $\phi$, the above three equations are actually two independent ones for the two unknowns $a(t)$ and $\phi(t)$.
The method of slow-roll parameters is usually used to obtain the equations of motion in the inflation. In this case, since action (5) has an additional term including a scalar field dependence, we are required to define new slow-roll parameters in addition to the usual slow-roll parameters used with the Einstein-Hilbert action. Such definitions should be in a way to ensure that the slow-roll conditions are still satisfied and the derivatives of $\xi(\phi)$ varies slowly. According to this requirement, the slow-roll parameters can be introduced as [47]
$\epsilon_{1}=-\dfrac{\dot{H}}{H^{2}}=-\dfrac{d\ln H}{d\ln a},\qquad\epsilon_{n+1}=\dfrac{d\ln\mid\epsilon_{n}\mid}{d\ln a},$ (14)
$\delta_{1}=4H\dot{\xi},\qquad\delta_{n+1}=\dfrac{d\ln\mid\delta_{n}\mid}{d\ln a},$ (15)
for $n\geqslant 1$ as an integer number, and where at least [^2] Otherwise, the coupling function $\xi$ is constant, and hence, there will be no effect of the Gauss-Bonnet term in four-dimensional spacetime. $\dot{\xi}\neq 0$. Thereupon, the approximation of the slow-roll parameters in the inflation are $\mid\epsilon_{n}\mid\ll 1$ and $\mid\delta_{n}\mid\ll 1$. By imposing these conditions, one obtains
$\dot{\phi}^{2}\ll V(\phi),\qquad\quad\mid\ddot{\phi}\mid\ll\,12H^{4}\mid\!\xi^{{}^{\prime}}\!\mid,$ (16)
$4H\mid\dot{\xi}\mid\ll 1,\qquad\quad\mid\ddot{\xi}\mid\ll\,H\mid\dot{\xi}\mid.$ (17)
Accordingly, the equations of motion (11), (12) and (13) can be rewritten as
$H^{2}\simeq\dfrac{1}{3}V,$ (18)
$\dot{H}\simeq-\dfrac{1}{2}\dot{\phi}^{2}-2H^{3}\dot{\xi},$ (19)
$\dot{\phi}\simeq-\dfrac{1}{3H}\left(V^{{}^{\prime}}+12H^{4}\xi^{{}^{\prime}}\right).$ (20)
Also, in this case, the $e$-folding number throughout the inflationary epoch generally is [59, 60]
$N(\phi)=\int^{\phi_{\rm end}}_{\phi}\dfrac{H}{\dot{\phi}}d\phi\simeq\int_{\phi_{\rm end}}^{\phi}\dfrac{V}{V^{{}^{\prime}}+4\xi^{{}^{\prime}}V^{2}/3}d\phi,$ (21)
where we have utilized the approximate Eqs. (18)-(20), and the subscript ‘end’ denotes the value of the inflaton scalar field at the end of inflation, i.e., when $\epsilon_{1}\!\!\mid_{\phi_{\rm end}}\simeq 1$.
Now, to facilitate the writing of equations, the quantity $Q(\phi)\equiv V^{{}^{\prime}}/V+4\xi^{{}^{\prime}}V/3$ can be introduced, with which the $e$-folding number and the slow-roll parameters can be rewritten as follows while using the approximate Eqs. (18)-(20).
$N(\phi)\simeq\int_{\phi_{\rm end}}^{\phi}\dfrac{d\phi}{Q},$ (22)
$\epsilon_{1}\simeq\frac{Q}{2}\frac{V^{{}^{\prime}}}{V},$ (23)
$\epsilon_{2}\simeq-Q\left(\frac{V^{{}^{\prime\prime}}}{V^{{}^{\prime}}}-\frac{V^{{}^{\prime}}}{V}+\frac{Q^{{}^{\prime}}}{Q}\right),$ (24)
$\delta_{1}\simeq-\frac{4Q}{3}\xi^{{}^{\prime}}V,$ (25)
$\delta_{2}\simeq-Q\left(\frac{\xi^{{}^{\prime\prime}}}{\xi^{{}^{\prime}}}+\frac{V^{{}^{\prime}}}{V}+\frac{Q^{{}^{\prime}}}{Q}\right).$ (26)
Also for later use, let us recall the following quantities. In Refs. [8, 24, 33, 35], the primordial power spectrum of scalar (curvature) and tensor perturbations is derived to be
$\mathcal{P}_{\rm S}=\frac{H^{2}}{4\pi^{2}c_{\rm S}^{3}F_{\rm S}},$ (27)
$\mathcal{P}_{\rm T}=\frac{2H^{2}}{\pi^{2}c_{\rm T}^{3}F_{\rm T}},$ (28)
respectively for the time of horizon crossing at $c_{\rm S}k=aH$ and $c_{\rm T}k=aH$ with the comoving wavenumber $k$, where to the lowest-order in the slow-roll parameters the difference of time of horizon crossing is unimportant [8]. Also, the time derivative of the flow parameters is neglected during the slow-roll inflation, and the squared propagation speed of the scalar and tensor perturbation modes and the other quantities appearing in relations (27) and (28) are [61, 62]
$c_{\rm S}^{2}=1+\frac{8\dot{\xi}\left[4\dot{H}\dot{\xi}+\Delta H(\ddot{\xi}-H\dot{\xi})\right]}{F_{\rm T}F_{\rm S}(1-\Delta/2)^{2}},$ (29)
$c_{\rm T}^{2}=1-\frac{4(\ddot{\xi}-H\dot{\xi})}{F_{\rm T}},$ (30)
where
$F_{\rm S}\equiv\frac{\dot{\phi}^{2}+6\Delta H^{3}\dot{\xi}}{(1-\Delta/2)^{2}H^{2}},$ (31)
$F_{\rm T}\equiv 1-4H\dot{\xi},$ (32)
and $\Delta\equiv(1-F_{\rm T})/F_{\rm T}$. Accordingly, the tensor-to-scalar ratio $r\equiv\mathcal{P}_{\rm T}/\mathcal{P}_{\rm S}$ and the scalar and the tensor spectral indices can be obtained in terms of the slow-roll parameters as
$r\simeq 8\mid\!2\epsilon_{1}-\delta_{1}\!\mid,$ (33)
$n_{\rm S}-1\simeq-2\epsilon_{1}-\frac{2\epsilon_{1}\epsilon_{2}-\delta_{1}\delta_{2}}{2\epsilon_{1}-\delta_{1}},$ (34)
$n_{\rm T}\simeq-2\epsilon_{1}.$ (35)
3 The Proposed Model
We intend to investigate the above results for a model where the inflaton scalar field is a Higgs scalar field. In this model, the potential can be the Higgs field potential, i.e., $V(\phi)=V_{0}(\phi^{2}-\nu_{\phi}^{2})^{2}$. However, since the inflaton scalar field in inflation is much larger than the vacuum expectation value of the Higgs scalar field, i.e., $\phi\gg\nu_{\phi}$, the potential can be approximated as the quartic form [^3] The quartic self-interaction coefficient $V_{0}$ is determined by the experimentally measured values of the mass of Higgs boson and its vacuum expectation value. In this case, the value of $V_{0}$ is approximately of order ${\cal O}(10^{-1})$, see, e.g., Ref. [63]. However, if needed as an inflaton scalar field, its value can be approximately of order ${\cal O}(10^{-13})$, see, e.g., Ref. [2].
$V(\phi)\simeq V_{0}\phi^{4},$ (36)
where $V_{0}$ is a non-zero constant. Moreover, in this model, we consider the Gauss-Bonnet coupling function $\xi(\phi)$ to be a particular choice, actually a dilaton-like coupling function like
$\xi(\phi)=\xi_{0}e^{-\lambda\phi},$ (37)
where $\xi_{0}$ and $\lambda$ are non-zero constants, $\phi$ is not constant, and in fact $\dot{\phi}\neq 0$. Although, for convenience, one can reasonably reduce the number of free parameters by introducing a parameter as $\alpha\equiv 4V_{0}\xi_{0}/3$.
Furthermore, in this model, the slow-roll parameters (23)-(26) are now functions of the inflaton scalar field as
$\epsilon_{1}\simeq\frac{2(4-\alpha\lambda e^{-\lambda\phi}\phi^{5})}{\phi^{2}},$ (38)
$\epsilon_{2}\simeq\frac{8-\alpha\lambda e^{-\lambda\phi}\phi^{5}(\lambda\phi-3)}{\phi^{2}},$ (39)
$\delta_{1}\simeq\frac{\alpha\lambda e^{-\lambda\phi}\phi^{5}(4-\alpha\lambda e^{-\lambda\phi}\phi^{5})}{\phi^{2}},$ (40)
$\delta_{2}\simeq\frac{4(\lambda\phi-3)-2\alpha\lambda e^{-\lambda\phi}\phi^{5}(\lambda\phi-4)}{\phi^{2}}.$ (41)
Figure 1: $y(\phi)=\alpha\lambda e^{-\lambda\phi}\phi^{5}/4$ versus $\phi$ for different values of $\alpha$ and $\lambda$ while satisfying condition (51).
Figure 2: $y(\phi)=\alpha\lambda e^{-\lambda\phi}\phi^{5}/4$ versus $\phi$ for a fixed value of $\alpha=0.01$ and different values of $\lambda$.
In turn, the tensor-to-scalar ratio and the scalar spectral index (33) and (34) can also be obtained as a function of the inflaton scalar field, i.e.,
$r\simeq\frac{8(4-\alpha\lambda e^{-\lambda\phi}\phi^{5})^{2}}{\phi^{2}},$ (42)
$n_{\rm S}-1\simeq\frac{-24+\alpha\lambda e^{-\lambda\phi}\phi^{5}(2\lambda\phi-4)}{\phi^{2}}.$ (43)
Before proceeding, investigating the proposed model in the absence of the Gauss-Bonnet correction is also instructive. That is, considering the general relativity limit with the quartic potential when $\xi_{0}=0$ (that results in $\alpha=0$). In this respect, the slow-roll parameters (38) and (39) reduce to the relevant ones in general relativity, [^4] In this case, the other two slow-roll parameters (40)and (41) that were introduced due to the presence of the Gauss-Bonnet term coupled to the inflaton scalar field, no longer exist, i.e., by relations (15), $\delta_{n}=0$ for all $n$. i.e.,
$\epsilon_{1}\simeq\frac{8}{\phi^{2}}\simeq\epsilon_{2}.$ (44)
In addition, the tensor-to-scalar ratio and the scalar spectral index (42) and (43) are
$r\simeq\frac{128}{\phi^{2}}\qquad{\rm and}\qquad n_{\rm S}-1\simeq\frac{-24}{\phi^{2}}.$ (45)
Utilizing relation (21), while using the value of $\phi^{2}_{\rm end}\simeq 8$ from relation (44) in the absence of the Gauss-Bonnet correction, the relation $\phi^{2}(N)\simeq 8(N-1)$ is obtained, and hence relations (45) read
$r\simeq\frac{16}{N-1}\qquad{\rm and}\qquad n_{\rm S}\simeq\frac{N-4}{N-1}.$ (46)
Thus, in this case, when we choose $N=70$, the best values for inflationary observables can be achieved as
$r\simeq 0.2319\qquad{\rm and}\qquad n_{\rm S}\simeq 0.9565,$ (47)
which roughly recover the predictions of the chaotic inflation model [59]. However, in the case of absence of the Gauss-Bonnet correction, the value obtained above for $r$ is not consistent with the Planck observable data.
Now, in the presence of the Gauss-Bonnet term, to proceed and compare the obtained results with the Planck observational data, the functionality of $\phi$ should be specified and then inserted into relations (42) and (43). In this regard, a routine procedure is to first acquire the inflaton scalar field as a function of the $e$-folding number. Thereupon, by inserting the best value of $e$-folding number (around 607060-70), the obtained results can be checked easily. Hence, we must solve relation (22) by knowing the inflaton scalar field at the end of inflation and then obtain the $e$-folding number as a function of the inflaton scalar field. Next, by deriving the inverse function, we can achieve the inflaton scalar field as a function of the $e$-folding number.
In this regard, for the assumed potential and the coupling function, relation (22) is now
$N(\phi)\simeq\int^{\phi}_{\phi_{\rm end}}\frac{\phi d\phi}{4-\alpha\lambda e^{-\lambda\phi}\phi^{5}}.$ (48)
However, even with the slow-roll approximations holding true, this integral is complicated and cannot be solved analytically. Hence, to facilitate the study and solvability of the system, certain approximations should be applied. In this respect, to proceed, we first utilize the well-known Taylor expansion
$\frac{1}{1-x}=\sum_{n=0}^{\infty}x^{n}\qquad{\rm for\!:}\ \mid x\mid<1,$ (49)
upon which, relation (48) reads
$N(\phi)\simeq\sum_{n=0}^{\infty}4^{-n-1}(\alpha\lambda)^{n}\int^{\phi}_{\phi_{\rm end}}e^{-n\lambda\phi}\phi^{5n+1}d\phi,$ (50)
when
$\vline\ \frac{\alpha\lambda}{4}e^{-\lambda\phi}\phi^{5}\ \vline<1.$ (51)
Then, within this condition, by analytically solving relation (50), we obtain
$N(\phi)\simeq\sum_{n=0}^{\infty}4^{(-n-1)}(\alpha\lambda)^{n}(n\lambda)^{(-5n-2)}\Big[\gamma(5n+2;n\lambda\phi)-\!\gamma(5n+2;n\lambda\phi_{\rm end})\Big],$ (52)
Figure 3: $N(\phi)$ versus $\phi$ for a fixed value of $\alpha=0.01$ and different values of $\lambda$.
Figure 4: $N(\phi)$ versus $\phi$ for different values of $\alpha$ with the corresponding minimum value of $\lambda$ in each case.
where $\gamma$ is the incomplete gamma function.
Before proceeding, let us examine condition (51) carefully. The functionality of $y(\phi)\equiv\alpha\lambda e^{-\lambda\phi}\phi^{5}/4$ indicates that its value is always positive and has a maximum for every positive value of $\alpha$ and $\lambda$, for instance, see Fig. (2). The value of $\phi$ for the peak of the curve of $y(\phi)$ is $\phi|_{{}_{\rm P}}=5/\lambda$, and hence $y_{\rm max}(\phi|_{{}_{\rm P}})=e^{-5}5^{5}\alpha\lambda^{-4}/4$. Thus, by condition (51), we obtain the constraint
$\alpha\lambda^{-4}<4\,e^{5}5^{-5}$ (53)
on positive values of $\alpha$ and $\lambda$. Alternatively, for any fixed value of $\alpha$, there is a lower-bound for $\lambda$ that satisfies condition (51), for instance, see Fig. (2) for $\alpha=0.01$ where $\lambda$ can have any value in the interval $0.48,\infty)$. Moreover, Fig. (2) illustrates that with a fixed value of $\alpha$, any increase in the value of $\lambda$ from its lower-bound makes a lower value for $y(\phi)$. Also, it would be constructive to have the slow-roll parameters (38)-(41) and the inflationary observables (42) and (43) in terms of the function $y(\phi)$ as follows.
$\epsilon_{1}\simeq 8(1-y)/\phi^{2},$ (54)
$\delta_{1}\simeq 16y(1-y)/\phi^{2},$ (55)
$\epsilon_{2}\simeq 4\left[2-y(\lambda\phi-3)\right]/\phi^{2},$ (56)
$\delta_{2}\simeq 4\left[\lambda\phi-3-2y(\lambda\phi-4)\right]/\phi^{2},$ (57)
$r\simeq 128(1-y)^{2}/\phi^{2},$ (58)
$n_{\rm S}-1\simeq 4\left[-6+y(2\lambda\phi-4)\right]/\phi^{2},$ (59)
where, by inverting the defined function $y(\phi)$, $\phi(y)=-5\,W[-\lambda(4y/\alpha\lambda)^{1/5}/5]/\lambda$, where $W$ is the Lambert or the product logarithm function.
On the other hand, since the main purpose of this work is to implement and investigate the impact of the proposed modified gravity with the specific model on the inflationary observables, to proceed, it is first required to keep the model parameters such that those lead to acceptable values for the $e$-folding number. Hence, to incorporate such a constraint and to ensure the correctness of the values used for the model parameters, we have also drawn the $e$-folding function (52) versus $\phi$ for different values of the model parameters in figures (4) and (4). Indeed, we probed to find acceptable values for each of $\alpha$ and $\lambda$ such that the values of $N$ fall within the acceptable range of observational data. However, in drawing the figures, the values chosen for $\alpha$ and $\lambda$ must satisfy constraint (53). In this regard, Fig. (4) shows that with a fixed value of $\alpha$, there is also an upper-bound for the corresponding $\lambda$ where acceptable values for the $e$-folding number can be achieved, e.g., in this figure, with $\alpha=0.01$, $\lambda$ has a suitable range $[0.48,0.51)$. Also, with similar check, Fig. (4) illustrates that there is an upper-bound for $\alpha$. Note that, in this figure, the plots are drawn for each $\alpha$ with the corresponding minimum value of $\lambda$ specified by constraint (53). In general, our investigations show that to have acceptable values for the $e$-folding number, $\alpha$ must be less than $0.2$ and $\lambda$ has a lower-bound specified by constraint (53) and an upper-bound such that relation (52) is satisfied.
Figure 5: $\phi(N)$ versus $N$ for fixed value of $\alpha=0.01$ and $\lambda=0.48$. For values of NN between around $30$ to $80$, the best curve fits analytically.
Now let us return to the main issue of attaining the values of $r$ and $n_{\rm S}$, which are essentially obtained by determining the function $\phi$ from the inversion of relation (52). Of course, first we must obtain the inflaton scalar field at the end of inflation from relation (38). Once again, even with the slow-roll approximations holding true, the system of most equations still remains intricate and cannot be solved analytically, and further considerations are needed to elucidate the inflationary phenomenology. In this regard, due to the analytical complexities of relation (38), we were forced to proceed for specified values of $\alpha$ and $\lambda$, while utilizing the Taylor expansion for the exponential function in that relation. Thus, first, from relation (38), we numerically obtain the inflaton scalar field at the end of inflation for $\alpha=0.01$ and $\lambda=0.48$ to be approximately equal to $2.75565$ and for $\alpha=0.02$ and $\lambda=0.57$ to be approximately equal to $2.70126$. Nevertheless, relation (52) is also intricate and cannot be inverted analytically, and numerical analysis methods must also be used. Our approach is to plot $N(\phi)$ versus $\phi$ with some appropriate values of $\alpha$ and $\lambda$, and then invert this plot, see, e.g., Fig. (5). Thereupon, while using numerical analysis methods, we have scrutinized to find the best fit to such original curve $\phi(N)$ at least in the range of $N$ values between around $30$ to $80$. For instance, Fig. (5) is the inversion of Fig. (4) for $\alpha=0.01$ and $\lambda=0.48$, wherein the best fitted curve (and its corresponding equation) has also been specified at least in the mentioned range for $N$.
At last, with some investigations, we specify a few best fitted functions for $\phi$ in terms of the $e$-folding number as follows. For instance, for $\alpha=0.01$ with $\lambda=0.48$ and $\lambda=0.49$ while using the related inflaton scalar field at the end of inflation in each case, the corresponding fitted functions are, respectively, the linear functions
$\phi(N)\simeq 0.0197N+8.4524,$ (60)
$\phi(N)\simeq 0.0404N+8.2923.$ (61)
Also for $\alpha=0.02$ with $\lambda=0.57$ and $\lambda=0.58$ while again using the corresponding $\phi_{\rm end}$ in each case, we obtain, respectively, the linear functions
$\phi(N)\simeq 0.0173N+7.3263,$ (62)
$\phi(N)\simeq 0.0440N+6.8702.$ (63)
Note that, these fitted linear functions are only suitable during inflation for the range of $N$ values between around $30$ to $80$ and obviously do not have graceful exit. Furthermore, over a wider range of the model parameters, no appropriate fitted function can be obtained, not even a linear one. Also, for some values of $\lambda$, a corresponding fitted curve takes on a more complicated shape. Nonetheless, from the point of view of numerical analysis, the important point is that the obtained fitted curves can still be applied during inflation.
Now, by substituting the fitted Eqs. (60)-(63) into relations (42) and (43), we obtain the relations of $r$ and $n_{\rm S}$ with respect to the $e$-folding number and the model parameters. Then, to get the best results comparable with the data (1) and (2), using the obtained relations, we have drawn $r$ versus $n_{\rm S}$ in Fig. 6 for a few acceptable $e$-folding numbers.
Figure 6: The tensor-to-scalar ratio $r$ versus the scalar spectral index $n_{\rm S}$ for two values of $\alpha$ each with corresponding two values of $\lambda$.
This figure shows that for two sets of the model parameters, i.e., $\alpha=0.01$ with $\lambda=0.48$ and $\alpha=0.02$ with $\lambda=0.57$, the obtained results for the inflationary observables are in good agreement with the observational data. Our investigations show that to have acceptable values for the inflationary observables consistent with the Planck observational data, the best model parameters in the range $N\in(50,70]$ should be in the range $\alpha\in(0,0.02]$ and slightly more with (almost) the corresponding minimum value of $\lambda$ specified by constraint (53) for each $\alpha$. The observable quantities are sensitive to the parameter $\alpha$, and when $\alpha$ increases from $0.02$, the results begin to deviate from the Planck observational data, and thus there is a small range for values of the model parameters to provide acceptable results.
For the above mentioned two optimal sets of the model parameters, let us record the obtained values, for instance, with $N=70$, by inserting Eqs. (60) and (62) into relations (42) and (43). Accordingly, the obtained results are, respectively,
$r\simeq 0.0359\times 10^{-2}\qquad{\rm and}\qquad 0.0035\times 10^{-2},$ (64)
$n_{\rm S}\simeq 0.9730\qquad{\rm and}\qquad 0.9839.{}\hskip 28.45274pt{}$ (65)
Also, the obtained values for the slow-roll parameters, relations (38)-(41), during inflation are, respectively,
$\epsilon_{1}\simeq 0.136\times 10^{-2}\qquad\,\ {\rm and}\qquad 0.049\times 10^{-2},$ (66)
$\epsilon_{2}\simeq 1.280\times 10^{-2}\qquad\,\ {\rm and}\qquad 0.780\times 10^{-2},$ (67)
$\delta_{1}\simeq 0.2680\times 10^{-2}\qquad{\rm and}\qquad 0.09754\times 10^{-2},$ (68)
$\delta_{2}\simeq 1.261\times 10^{-2}\qquad\,\ {\rm and}\qquad 0.7764\times 10^{-2}$ (69)
that are well below one as expected. [^5] Since we have used numerical analysis methods, one needs to ensure these smallnesses. In turn, utilizing relation (35), the tensor spectral index is obtained, respectively,
$n_{\rm T}\simeq-0.272\times 10^{-2}\qquad{\rm and}\qquad-0.098\times 10^{-2}.$ (70)
Moreover, the value of $y(\phi)$ function is approximately $0.9835$ and $0.9955$, respectively, that is less than one as expected.
It is noteworthy that the sufficiently low value obtained for the tensor-to-scalar ratio is comfortably below the experimental limits in the primordial gravitational waves. However, the value obtained for it in the absence of the Gauss-Bonnet term, relation (47), is far from the Planck observable data. Thus, its new consistent values represent the significant effect of including the Gauss-Bonnet correction. It is worth noting that, utilizing the values of the mentioned two optimal sets of the model parameters and $N=70$ as well as Eqs. (60) and (62) to estimate the two values of $V_{0}$ mentioned in the Footnote $3$, the estimated value of the coefficient of the Gauss-Bonnet term in action (5) (i.e., $\xi(\phi)/2$) is approximately of order ${\cal O}(10^{-5})$ or ${\cal O}(10^{9})$, respectively.
At this stage, let us examine the speed of gravitational waves during inflation. In this regard, relation (30) shows the propagation speed of tensor perturbation modes (including gravitational waves) in the presence of the Gauss-Bonnet correction. During inflation, utilizing the approximations in relation (17) into relation (30), we obtain
$c_{{}_{\rm GW}}^{2}\!\simeq\!1\!+\frac{4H\dot{\xi}}{1\!-\!4H\dot{\xi}}\!\simeq\!1\!+4H\dot{\xi}(1\!+4H\dot{\xi})\!\simeq\!1\!+4H\dot{\xi}\!=\!1\!+\delta_{1},$ (71)
where in the last equality we have used relation (15). Thus, under the approximate conditions used for inflation, if the time derivative of the coupling function is positive, the speed of gravitational waves during inflation will be greater than the speed of light. Now, using relation (40), while writing $\phi(N)$, relation (71) gives the speed of gravitational waves during inflation in terms of the $e$-folding as
$c_{{}_{\rm GW}}^{2}(N)\!\simeq\!1\!+\alpha\lambda e^{-\lambda\phi(N)}\phi^{3}(N)\left[4\!-\!\alpha\lambda e^{-\lambda\phi(N)}\phi^{5}(N)\right].$ (72)
Then, utilizing the linear function of the fitted scalar field, Eqs. (60)-(63), and using the values of the two optimal sets of the model parameters $\alpha$ and $\lambda$ obtained in the preceding section, we have plotted the speed of gravitational waves within inflation in Fig. (7).
Figure 7: The speed of gravitational waves $c_{{}_{\rm GW}}$ versus $N$ for two optimal pairs of $\alpha$ and $\lambda$ in the range of $N$ from $30$ to $80$ during inflation.
This figure shows that the speed of gravitational waves during inflation is greater than one (i.e., the speed of light) but very close to it, and its very slight difference from one is due to the presence of the Gauss-Bonnet correction. Hence, the contribution of the Gauss-Bonnet term to inflationary spectra in this Higgs inflation-like model during inflation is almost negligible, although not absolutely zero. Of course, if $\dot{\xi}=0$ in relation (71), this speed will be exactly equal to one and the Gauss-Bonnet correction has no effect at all. In fact, in such a situation, since $\xi$ will be a constant value, the Gauss-Bonnet term will have no effect on the field equations in four dimensions.
Furthermore, at the end of inflation, utilizing $\epsilon_{1}\!\!\mid_{\phi_{\rm end}}\simeq 1$ from relation (38) in relation (40), we arrive at
$\delta_{1}\!\!\mid_{\phi_{\rm end}}\simeq 2-\frac{\phi_{\rm end}^{2}}{4}.$ (73)
Then, by substituting it into relation (71), we obtain the speed of gravitational waves at the end of inflation as
$c_{{}_{\rm GW}}^{2}\!\!\mid_{\phi_{\rm end}}\simeq 3-\frac{\phi_{\rm end}^{2}}{4}.$ (74)
In Table (1), we have presented its values for the values of the two optimal sets of the model parameters $\alpha$ and $\lambda$.
Table 1: The speed of gravitational waves at the end of inflation for the values of $\alpha$ and $\lambda$.
$\alpha=$ $\lambda=$ $\phi_{\rm end}\simeq$ $c_{{}_{\rm GW}}\!\!\mid_{{}_{\phi_{\rm end}}}\simeq$
0.01 0.48 2.7557 1.0496
0.49 2.7561 1.0493
0.02 0.57 2.7013 1.0843
0.58 2.7023 1.0837
Although the Gauss-Bonnet term exerts a significant influence on the underlying physics of the system at the end of the inflationary era, its very slight difference from one justifies that if gravitons are massless after the inflationary era, compatibility can be achieved by equating the speed of gravitational waves to the speed of light or infinitesimally close to it.
4 Conclusions
We have considered the Gauss-Bonnet term coupled with the Higgs scalar field added to the Einstein-Hilbert action. Then, we have derived the equations of motion and, while employing the FLRW metric, the modified Friedmann equations. Thereafter, we have approximated these equations during inflation by considering the appropriate slow-roll parameters. Moreover, the proposed model includes the Higgs potential and the coupling coefficient as a dilaton-like coupling.
Under these assumptions, we encountered calculations to obtain the tensor-to-scalar ratio and the scalar spectral index that are rather complicated and cannot be easily solved analytically. Hence, further approximations have been needed in order to derive the inflationary phenomenology. In this regard, in the first stage, we did not proceed directly with numerical analysis methods, and instead we have found that it is productive by first employing a well-known Taylor expansion, we can then solve the $e$-folding integral analytically. Meanwhile, this procedure has caused a constraint on the model parameters. However, we still need to determine the function of $e$-folding through numerical analysis methods for a certain range of derived values for the model parameters in order to have acceptable values for it.
In the process of this work, several plots have been drawn and at the same time numerical analysis methods have been used. Thereupon, in the specified range of the model parameters, we have drawn the inverse function of the $e$-folding function numerically and scrutinized to find the best fitted curve for the inflaton scalar field in an acceptable range of the $e$-folding number. At last, while utilizing the best fitted function for the inflaton scalar field, in order to attain acceptable values for the tensor-to-scalar ratio and the scalar spectral index, we have plotted them against each other with the appropriate values obtained for the model parameters such that the $e$-folding values fall within the acceptable range of observational data.
Finally, for the proposed model, with a certain range of values derived for the model parameters, we have obtained results for the tensor-to-scalar ratio and the scalar spectral index that are in good agreement with observational data within the acceptable range of the $e$-folding values. However, there is not much choice for the values of the model parameters. Of course, one of the key points of this work is to achieve results through analytical analysis methods. Also, in the absence of the Gauss-Bonnet term, we find that the inflationary observables are roughly the same as the predictions of the chaotic inflation model. However, in this case, the value obtained for the tensor-to-scalar ratio is not consistent with the Planck observable data, while the very acceptable values obtained for it in the presence of the Gauss-Bonnet term indicate the significant effect of including the Gauss-Bonnet correction. In addition, the results of the model show that the speed of gravitational waves during inflation is greater than one (i.e., the speed of light) but very close to it, and its very slight difference from one is due to the presence of the Gauss-Bonnet correction.
Acknowledgments
The authors thank the Deputy for Research and Technology of Shahid Beheshti University.